# Identification and estimation of vaccine effectiveness in the test-negative design under equi-confounding

### v3

@@ Appendix D Doubly robust estimator under cohort sampling
[…]
@@
D.1 Proof of equivalence with UDiD
@@
[…]
@@
D.2 Proof of double robustness
@@ D.3 Estimation
Appendix E Doubly robust estimator under TND sampling
@@
[…]
@@
E.1 Proof of double robustness
@@ E.2 Estimation
Appendix F Direct effects of vaccination and identification of alternative estimands
@@ F.1 Symptomatic infection
$=\dfrac{\Pr(I^{1}=2|T=1,V=1)}{E\left\{\Pr(I^{0}=2|T=1,V=1,X)\Big{|}T=1,V=1\right\}}$
$=\dfrac{\Pr(I^{1}=2|T=1,V=1)}{E\left\{\dfrac{\Pr(I^{0}=2|V=1,X)}{\Pr(I^{1}=2|V=1,X)}\Pr(I^{1}=2|T=1,V=1,X)\Big{|}T=1,V=1\right\}}$
$=\dfrac{\Pr(I^{1}=2|T=1,V=1)}{E\left\{\dfrac{\Pr(I^{0}=1|V=1,X)}{\Pr(I^{0}=1|V=0,X)}\dfrac{\Pr(I^{0}=2|V=0,X)}{\Pr(I^{1}=2|V=1,X)}\Pr(I^{1}=2|T=1,V=1,X)\Big{|}T=1,V=1\right\}}$
$=\dfrac{\Pr(I^{1}=2|T=1,V=1)}{E\left\{\dfrac{\Pr(I^{1}=1|V=1,X)}{\Pr(I^{0}=1|V=0,X)}\dfrac{\Pr(I^{0}=2|V=0,X)}{\Pr(I^{1}=2|V=1,X)}\Pr(I^{1}=2|T=1,V=1,X)\Big{|}T=1,V=1\right\}}$
$=\dfrac{\Pr(I=2|T=1,V=1)}{E\left\{\dfrac{\Pr(I=1|V=1,X)}{\Pr(I=1|V=0,X)}\dfrac{\Pr(I=2|V=0,X)}{\Pr(I=2|V=1,X)}\Pr(I=2|T=1,V=1,X)\Big{|}T=1,V=1\right\}}$
$=\dfrac{\Pr(I=2|T=1,V=1)}{E\left\{\dfrac{\Pr(I=1|T=1,V=1,X)}{\Pr(I=1|T=1,V=0,X)}\dfrac{\Pr(I=2|T=1,V=0,X)}{\Pr(I=2|T=1,V=1,X)}\Pr(I=2|T=1,V=1,X)\Big{|}T=1,V=1\right\}}$
$=\dfrac{\Pr(I=2|T=1,V=1)}{E\left\{\dfrac{\Pr(I=1|T=1,V=1,X)}{\Pr(I=1|T=1,V=0,X)}\Pr(I=2|T=1,V=0,X)\Big{|}T=1,V=1\right\}}$
$=\dfrac{\Pr(Y=2|S=1,V=1)}{E\left\{\Pr(Y=2|S=1,V=0,X)\dfrac{\Pr(Y=1|S=1,V=1,X)}{\Pr(Y=1|S=1,V=0,X)}\Big{|}S=1,V=1\right\}}$
@@
Appendix G Extended discussion of equi-confounding assumption
G.1 Unmeasured health-seeking behavior
Among the stated motivations for the test-negative design when it was first introduced was the ability of the TND to adjust for unmeasured health-seeking behavior. Early observational studies of the effectiveness of flu vaccines tended to overstate their effectiveness relative to randomized trials. At the time it was hypothesized that the bias may have been caused by differential health-seeking behavior. Because testing was not mandatory and only available at health facilities, in order to be a confirmed case individuals had to be motivated to seek care and it was assumed that the individuals who did this were both healthier and more likely to be vaccinated than the general population. By limiting to those receiving a test, it was claimed that the TND resolved this issue, at least in the case that health seeking could be considered a binary construct. However, subsequent studies pointed out that doing so risked selection bias.
Here we show how unmeasured health-seeking or, test-seeking as we characterize it, fits into the equi-confounding framework. Consider the equi-selection assumption (A3b)
$\dfrac{\Pr(Y^{0}=1|V=1,X)}{\Pr(Y^{0}=2|V=0,X)}=\dfrac{\Pr(T^{0}=1|I^{0}=1,V=1,X)}{\Pr(T^{0}=1|I^{0}=1,V=0,X)}$
It captures the potential difference in receiving a test when symptomatic between those
G.2 Immortal time among vaccinated
Concerns have previously been raised about the potential for design-induced biases in the test-negative design resulting from failure to explicitly emulate a randomized trial. Here we show that, when the degree of design-induced bias is similar for test-positive and test-negative illnesses, they tend to cancel each other out and produce no bias in the TND estimate of vaccine effectiveness when they are equivalent.
$\frac{\Pr(Y^{0}=2|V=1,X)}{\Pr(Y^{0}=2|V=0,X)}=\eta_{2}(X)$
for some $\eta_{2}(X)<1$ as the person time in the first 28 days after receiving a vaccine is differentially discarded among those who received a vaccine compared to those who did not. However, if this exclusion of person time among the vaccinated is applied independently of the test result, the bias also applies to the incidence of test-negative illness such that
$\frac{\Pr(Y^{0}=1|V=1,X)}{\Pr(Y^{0}=1|V=0,X)}=\eta_{1}(X)$
for some $\eta_{1}(X)<1$. In the special case that $\eta_{2}(X)=\eta_{1}(X)$, which could occur if the trends in incidence of test-positive and test-negative illness are approximately parallel [^1] Note: the absolute number of cases need not be equal, we just need the trends in the incidence to be parallel over the 28 day period where immortal time is accruing, we have equi-confounding and the design-induced bias is completely removed by $\Psi^{*}_{om}$ and $\Psi^{*}_{ipw}$.
G.3 Correlated vaccination history
Prior work on the TND suggests that the design could be biased when vaccines for the test negative illness exist in addition to the focal vaccine for the test positive pathogen and when vaccination behavior is correlated. For instance, denote by $V^{*}$ the vaccination against the test-negative pathogen and assume first $V^{*}$ is unavailable to the investigator. As shown in the DAG in Figure A1 this would also violate our identifiability assumptions because it would imply the existence of a confounder that does not affect the test positive and test negative illnesses equally, i.e.,
G.4 Home testing
$X$$V$$I=2$$T$$I=1$$U$$V^{*}$
(a) Correlated vaccination behavior
$X$$V$$I=2$$R$$T$$I=1$$U$
(b) At home testing
@@ Appendix H Identification when test is imperfect
In the main text, we assumed the availability of a perfect test such that $\widetilde{Y}_{i}=Y_{i}$ for all individuals where $\widetilde{Y}$ is the test result and $Y$ is the true value. In the absence of such a test, here we describe the potential for bias due to misclassification of case status which has also been described previously [sullivan_theoretical_2016]. Because most studies employ real-time reverse-transcription polymerase chain reaction (RT-PCR), false positives are unlikely as most RT-PCR tests have specificity approaching 100%. Therefore, misclassification of test-positives is probably rare, perhaps due to sample contamination or data entry errors. False negatives may be more likely as test sensitivity is generally lower due to the fact that (1) some of those infected with test-positive illness may not shed detectable virus or viral RNA; (2) some may seek care after shedding has ceased; or (3) sample quality may be compromised due to swab quality or inadequate storage. If sample collection is sufficient such that the source of the test-negative infection can be identified via RT-PCR, then sensitivity with respect to the test-positive illness can be improved by limiting to test-negative controls with an identified cause.
@@ Appendix I Simulation details
Likewise, violations of Assumption A2 are controlled by $\beta_{1V}$ and $\tau_{1V}$. When $\beta_{1V}\neq 0$ there is a direct effect of vaccination on test-negative illness $I=1$ and when $\tau_{1V}\neq 0$ there is a direct effect of vaccination on testing among those with test-negative illness. When this occurs, we can at best identify the ratio of vaccine effects on the two illness outcomes, i.e. $\exp(\beta_{2V}+\tau_{2V})/\exp(\beta_{1V}+\tau_{1V})$, provided there is no effect modification by $X$ or $U$. As we show in Appendix F and illustrate further via simulation below, under the special case of equal effects on testing, i.e. $\tau_{2V}=\tau_{1V}$, certain effects on symptomatic illness may still be recovered.
@@
where $\mu_{v}(X)=\Pr(Y=2|X,V=v)$ and $\mu_{v}(X,U)=\Pr(Y=2|X,U,V=v)$ The first estimator represents the typical case that the covariate $U$ is unmeasured, and the second represents the hypothetical case where $U$ is available, allowing for conditioning on a sufficient adjustment set. We consider both the typical case that the covariate $U$ is unmeasured, and the hypothetical case where $U$ is available, allowing for conditioning on a sufficient adjustment set. Finally, we include estimates from and $\widehat{\Psi}_{om}$ based on $\Psi_{om}$ from Proposition 1 and discussed further in section C that uses the full sample instead of restricting to those tested. Finally, in scenario 8, we demonstrate the double robustness property of $\widehat{\Psi}_{dr}^{*}$ by adjusting the data generation process for $V$ to be
@@
$\beta_{20}$2.4 -2.4 -2.4 -2.4 -2.4 -2.4 -2.4 -2.4
$\beta_{2V}$1 -1 -1 -1 -1 -1 -1 -0.25
$\beta_{2X}$0.625 -0.625 -0.625 -0.625 -0.625 -0.625 -0.625 -0.625
@@
Table A3: Results of simulation study: 1000 simulations of sample size $n=$15,000 across 9 scenarios with TND selection probabilities between 9% and 15%. Bias is estimated mean bias, SE is estimated monte carlo standard error, Coverage is estimated coverage probability of 95% CI
@@
Table A4: Robustness of estimators to misspecification of nuisance functions: 1000 simulations of sample size $n=$15,000 from modified processes based on scenario 8 in Table 3. Bias is estimated mean bias, SE is estimated monte carlo standard error, Coverage is estimated coverage probability of 95% CI

### v4

@@ Appendix E Doubly robust estimator under cohort sampling
and $\beta(1,X)=\beta(2,X)$ under equi-confounding. [20] also provide a doubly-robust estimator of $\psi$, given by
@@
E.1 Proof of equivalence with UDiD
@@
$\quad=E\left[\dfrac{\int\mathbbm{1}(Y=2)\exp\{\beta(y,X)\}f(y|V=0,X)dy}{\int\exp\{\beta(y,X)\}f(y|V=0,X)}\bigg{|}V=1\right]$
@@
E.2 Proof of double robustness
@@ E.3 Estimation
Appendix F Doubly robust estimator under TND sampling
@@
is the log of the conditional odds ratio function. Furthermore, as described in [30], a doubly-robust estimator of $\phi^{*}(X)$ may be obtained as the solution to the empirical analogue of the population moment equation
@@
F.1 Proof of double robustness
@@ F.2 Estimation
Appendix G Direct effects of vaccination and identification of alternative estimands
@@ G.1 Symptomatic infection
[…]
@@ Appendix H Identification when test is imperfect
In the main text, we assumed the availability of a perfect test such that $\widetilde{Y}_{i}=Y_{i}$ for all individuals where $\widetilde{Y}$ is the test result and $Y$ is the true value. In the absence of such a test, here we describe the potential for bias due to misclassification of case status which has also been described previously [11]. Because most studies employ real-time reverse-transcription polymerase chain reaction (RT-PCR), false positives are unlikely as most RT-PCR tests have specificity approaching 100%. Therefore, misclassification of test-positives is probably rare, perhaps due to sample contamination or data entry errors. False negatives may be more likely as test sensitivity is generally lower due to the fact that (1) some of those infected with test-positive illness may not shed detectable virus or viral RNA; (2) some may seek care after shedding has ceased; or (3) sample quality may be compromised due to swab quality or inadequate storage. If sample collection is sufficient such that the source of the test-negative infection can be identified via RT-PCR, then sensitivity with respect to the test-positive illness can be improved by limiting to test-negative controls with an identified cause.
@@ Appendix I Simulation details
Likewise, violations of Assumption A2 are controlled by $\beta_{1V}$ and $\tau_{1V}$. When $\beta_{1V}\neq 0$ there is a direct effect of vaccination on test-negative illness $I=1$ and when $\tau_{1V}\neq 0$ there is a direct effect of vaccination on testing among those with test-negative illness. When this occurs, we can at best identify the ratio of vaccine effects on the two illness outcomes, i.e. $\exp(\beta_{2V}+\tau_{2V})/\exp(\beta_{1V}+\tau_{1V})$, provided there is no effect modification by $X$ or $U$. As we show in Appendix G and illustrate further via simulation below, under the special case of equal effects on testing, i.e. $\tau_{2V}=\tau_{1V}$, certain effects on symptomatic illness may still be recovered.
@@
where $\mu_{v}(X)=\Pr(Y=2|X,V=v)$ and $\mu_{v}(X,U)=\Pr(Y=2|X,U,V=v)$ The first estimator represents the typical case that the covariate $U$ is unmeasured, and the second represents the hypothetical case where $U$ is available, allowing for conditioning on a sufficient adjustment set. We consider both the typical case that the covariate $U$ is unmeasured, and the hypothetical case where $U$ is available, allowing for conditioning on a sufficient adjustment set. Finally, we include estimates from and $\widehat{\Psi}_{om}$ based on $\Psi_{om}$ from Proposition 1 and discussed further in section D that uses the full sample instead of restricting to those tested. Finally, in scenario 8, we demonstrate the double robustness property of $\widehat{\Psi}_{dr}^{*}$ by adjusting the data generation process for $V$ to be
@@
$\beta_{20}$ -2.4 -2.4 -2.4 -2.4 -2.4 -2.4 -2.4 -2.4
$\beta_{2V}$ -1 -1 -1 -1 -1 -1 -1 -0.25
$\beta_{2X}$ -0.625 -0.625 -0.625 -0.625 -0.625 -0.625 -0.625 -0.625
@@
Table A3: Results of simulation study: 1000 simulations of sample size n=n=15,000 across 9 scenarios with TND selection probabilities between 9% and 15%. Bias is estimated mean bias, SE is estimated monte carlo standard error, Coverage is estimated coverage probability of 95% CI
@@
Table A4: Robustness of estimators to misspecification of nuisance functions: 1000 simulations of sample size n=n=15,000 from modified processes based on scenario 8 in Table 3. Bias is estimated mean bias, SE is estimated monte carlo standard error, Coverage is estimated coverage probability of 95% CI